# Provable Privacy with Non-Private Pre-Processing

### v1

@@ 4.1 Sensitivity of Common Pre-Processing Algorithms
[…]
@@ 4.2 Privacy Analysis for Pre-Processing Algorithms
Deduplication
Quantization
Mean imputation
PCA
${\mathcal{M}}_{G}$
${1.05\alpha\varepsilon^{2}}\left({1+p^{2}}\right)$
${1.05\alpha\varepsilon^{2}}\left({1+\eta^{2}p^{2}}\right)$
${1.05\alpha\varepsilon^{2}}\left({1+\frac{4p^{2}}{(n-p)^{2}}}\right)$
${1.05\alpha\varepsilon^{2}}\left({1+\frac{12.2^{2}}{(\delta_{\min}^{k})^{2}}}\right)$
${\mathcal{A}}_{\mathrm{GD}}$
${1.05\alpha\varepsilon^{2}}\left({4+p^{2}}\right)$
${1.05\alpha\varepsilon^{2}}\left({4+\eta^{2}p^{2}}\right)$
${4.2\alpha\varepsilon^{2}}\left({1+\frac{p^{2}}{(n-p)^{2}}}\right)$
${1.05\alpha\varepsilon^{2}}\left({4+\frac{12.2^{2}}{(\delta_{\min}^{k})^{2}}}\right)$
${\mathcal{M}}_{L}/{\mathcal{M}}_{E}$
$\varepsilon\left({1+p}\right)$
$\varepsilon\left({1+\eta p}\right)$
$\varepsilon\left({1+\frac{2p}{n-p}}\right)$
$\varepsilon\left({1+\frac{12.2}{\delta_{\min}^{k}}}\right)$
${\mathcal{A}}_{\mathrm{SGD-samp}}$
${1.05\alpha\varepsilon^{2}}\left({2\alpha+\frac{12.2^{2}}{(\delta_{\min}^{k})^{2}}}\right)$
${\mathcal{A}}_{\mathrm{SGD-iter}}$
${1.1\alpha\varepsilon^{2}}\left({1+\frac{p^{2}n\log(n-p)}{(n-p)\log n}}\right)$
${1.1\alpha\varepsilon^{2}}\left({1+\frac{\eta^{2}p^{2}n\log(n-p)}{(n-p)\log n}}\right)$
${1.1\alpha\varepsilon^{2}}\left({1+\frac{4p^{2}n\log(n-p)}{(n-p)^{3}\log n}}\right)$
Table 2: Overall privacy guarantees of pre-processed DP pipelines with $\alpha\geq 11$. Let $p$ represent the $L_{\infty}$ sensitivity of deduplication, quantization and mean imputation. We also assume the size of the dataset $n\geq 101$ for PCA, and the Lipschitz and smoothness parameters, along with global sensitivity, are set to $1$.
@@
Figure 2: Visualization of the overall privacy of pre-processed Gaussian mechanism with $1$-Lipschitz function $f$ whose global sensitivity is $\Delta_{f}=1$, and privacy parameters $\epsilon=1,\delta=0.001$.
@@
Table 2 shows that the pre-processing methods discussed in Section 4.1 typically lead to a minimal, constant-order increase in the privacy cost for common DP mechanisms, depending on the datasets in ${\mathcal{L}}$. Specifically, deduplication results in a constant increase in the privacy cost when the datasets in ${\mathcal{L}}$ do not contain large clusters, i.e. $\Delta_{\infty}=o(n)$, whereas quantization can handle larger clusters of size $\Delta_{\infty}=o(\nicefrac{{n}}{{\eta}})$. The privacy cost of mean imputation remains constant for datasets with few missing values ($p=o(n)$). For PCA, we only present the results for rank reduction and the results for dimension reduction follows similarly. The additional privacy cost remains small for low rank datasets with $\delta_{\min}$ bounded away from 0.
@@ 5.1 Limitations of conditional privacy guarantee
It is important to recognise that datasets exhibiting pathologucal characteristics, like those above, are usually not of practical interest in data analysis. For instance, applying mean imputation to a dataset where the number of missing points is nearly equal to the size of the dataset or implementing approximate deduplication that results in the removal of nearly all data are not considered sensible practice. A possible solution to this problem is to potentially refuse to provide an output when the input dataset is deemed pathological, provided that the decision to refuse is made in a manner that preserves privacy. This approach is adopted in the following section, where we employ the Propose-Test-Release framework proposed by Dwork and Lei (2009) to establish unconditional privacy guarantees over all possible datasets, at the expense of accuracy on datasets that are “pathological”.
@@ 5.2 Unconditional privacy guarantees via PTR
We present Algorithm 1 that applies PTR procedure to combine the non-private PCA for rank reduction with DP-GD. It exhibits unconditional privacy guarantee over all possible datasets (Theorem 7). While this technique is specifically described for combining PCA with DP-GD, it’s worth noting that the same approach can be applied to other pairings of Differentially Private mechanisms and pre-processing algorithms, though we do not explicitly detail each combination since their implementations follow the same basic principles.
@@
Input: Dataset $S$, estimated lower bound $\beta$, privacy parameters $\varepsilon,\delta$, Lipschitzness $L$
@@
For any $L$-Lipschitz and $\mu$-smooth loss function $\ell$, $\epsilon>0$ and $\delta\leq\exp\left({-1.05\varepsilon^{2}\left({1+\frac{12.2^{2}\mu^{2}}{L^{2}\beta^{2}}}\right)}\right)$, Algorithm 1 with privacy parameters $\epsilon,\delta$, and estimated lower bound $\beta$ is $\left({\widehat{\varepsilon}+\varepsilon,\delta}\right)$-DP on a dataset of size $n\geq 101$, where $\widehat{\varepsilon}=3\varepsilon\sqrt{1.05\left({1+\frac{12.2^{2}\mu^{2}}{L^{2}\beta^{2}}}\right)\log\frac{1}{\delta}}$.
@@
For $\delta,\varepsilon$ defined in Algorithm 1, $n\geq 101$, and any $\beta\leq\delta_{k}\left({S}\right)$, with probability at least $1-\gamma$,  Algorithm 1 outputs $\hat{\theta}$ such that the excess empirical risk
@@
where $\gamma=\frac{1}{2\delta}\exp\left({-\frac{(\delta_{k}(S)-\beta)n\varepsilon}{12.2}}\right),\Lambda=\sum_{i=k+1}^{d}\lambda_{i}(S)$ where the high probability is over the randomness in Step 1 and the expectation is over the randomness of Step 6.
The results follow a similar analysis as Song et al. (2021), which shows that the convergence bound of DP-GD scales with the rank of the dataset. However, when the dataset remains full rank but the first $k$ eigenvalues dominates the rest, e.g. $\Lambda=O(\sqrt{k})$, the convergence bound following Song et al. (2021) is $O(\sqrt{d})$. In contrast, Proposition 8 leads to a dimension-independent convergence bound of order $O(\sqrt{k})$, with a slight degradation in the privacy guarantee. Here, $\beta$ introduces the trade-off between privacy and utility. Large $\beta$ leads to tighter privacy guarantee (small effective $\varepsilon$ in Theorem 7) at the risk of worse utility (large $\gamma$ in Proposition 8).
@@
We conducted experiments on a synthetic approximately low rank dataset to corroborate our results in Proposition 8, and summarised the results in Figure 3. For each overall privacy parameter $\varepsilon$, we evaluated the excess empirical risk of DP-SGD with: a) non-private PCA with an adjusted privacy parameter from Table 2, b) DP-PCA, with half of the privacy budget to allocated to DP-PCA, and c) no pre-processing.
The dataset’s approximate low rankness, indicated by small yet positive eigenvalues, prevents logistic regression without pre-processing from leveraging dimension-independent optimization guarantees (Proposition 8). Non-private PCA, while incurring a minor constant-order privacy cost, effectively utilizes the data’s low rankness and offers significant benefits especially at smaller $\varepsilon$ values where optimization error is dominated by the first term in Equation 48. Conversely, DP-PCA exhibits poorer performance due to an inherent error of order $\Omega(d)$ (Liu et al., 2022).
@@ 6 Conclusion
7 Acknowledgement
We thank Francesco Pinto for the initial discussion of the project. We also thank Alexandru Ţifea, Jiduan Wu and Omri Ben-Dov for their useful feedback.
@@ 7 Acknowledgement
In this section, we present the proofs for the Theorems in Section 3.1. Section A.1 contains the proofs for the properties of SRDP. Section A.2 provides the exact SRDP parameters of common DP mechanisms.
A.1 Proof of properties of SRDP
See 1
@@ A.1 Proof of properties of SRDP
Lemma A (Additivity of Renyi divergence [van Erven and Harremoës, 2012]).
@@
Lemma B (Weak convexity of Renyi divergence, Lemma 25 in Feldman et al. [2018]).
@@ A.2 Proof of Theorem 1(Derivation of Table 1)
[…]
We note that Definition 6 is different from the usual definition of smoothness in the literature that requires $\left\lVert\nabla_{\theta}\ell(\theta,x)-\nabla_{\theta}\ell(\theta^{\prime},x)\right\rVert_{2}\leq\mu\left\lVert\theta-\theta^{\prime}\right\rVert_{2}$ for all $x\in{\mathcal{K}}$ and $\theta,\theta^{\prime}\in\Theta$ [Chaudhuri et al., 2011, Feldman et al., 2018]. However, many loss functions, including square loss and logistic loss, are smooth on common model classes based on Definition 6.
@@ A.2.1 Output Perturbation and Random Sampling methods
This section provides the detailed theorems and proofs for Smooth RDP parameters of output perturbation methods (Gaussian and Laplace mechanism) and random sampling methods (Exponential mechanism) in Theorem 9, 10 and 11 respectively, as stated in Table 1.
@@
Lemma C (Corrolary 3 in Mironov [2017]).
@@
Laplace mechanism ${\mathcal{M}}_{L}$ is $(\varepsilon,0)$-DP [Dwork et al., 2006]. By the equivalence of $(\varepsilon,0)$-DP and $(\infty,\varepsilon)$-RDP [Mironov, 2017], ${\mathcal{M}}_{L}$ is $(\infty,\varepsilon)$-RDP. This implies that ${\mathcal{M}}_{L}$ is $(\alpha,\varepsilon)$-RDP for any $1\leq\alpha\leq\infty$ by the monotonicity of Rényi divergence (D).
Lemma D (Monotonicity of Rényi divergence Mironov [2017]).
@@
The proof for RDP of Exponential mechanism is similar to that of Laplace mechanism. As the exponential mechanism ${\mathcal{M}}_{E}$ is $(\varepsilon,0)$-DP [Dwork et al., 2006]. By the equivalence of $(\varepsilon,0)$-DP and $(\infty,\varepsilon)$-RDP [Mironov, 2017], ${\mathcal{M}}_{E}$ is $(\infty,\varepsilon)$-RDP. By the monotonicity of Rényi divergence (D), ${\mathcal{M}}_{E}$ is also $(\alpha,\varepsilon)$-RDP for any $\alpha\geq 1$.

### v4

@@ 4.1 Sensitivity of Common Pre-Processing Algorithms
where $\delta_{\min}^{k}({\mathcal{L}})=\min_{S\in{\mathcal{L}}}\lambda_{k}(S)-\lambda_{k+1}(S)$ is the minimum gap between the $k^{\text{th}}$ and $(k+1)^{\text{th}}$ eigenvalue over any covariance matrix of $S\in{\mathcal{L}}$.
Standard Scaling
Scaling is one of the most common pre-processing methods. In Proposition 6, we provide the sensitivity results of standard scaling which scales each feature to have mean 0 and standard deviation 1, and min max scaling, which scales each feature to the interval between 0 and 1.
Proposition 6.
For a dataset collection ${\mathcal{L}}$, the $L_{\infty}$ sensitivity of standard scaling and min max scaling is the size of the datasets in ${\mathcal{L}}$, i.e. $n$. The $L_{2}$ sensitivity of standard scaling is
$\Delta_{2}=\frac{2}{\sigma_{\min}^{3}n}+\frac{2}{n\sigma_{\min}},$
where $\sigma_{\min}$ is the minimum standard deviation over datasets in ${\mathcal{L}}$.
@@ 4.2 Privacy Analysis for Pre-Processing Algorithms
Table 2: Overall privacy guarantees $\widehat{\varepsilon}$ of pre-processed DP pipelines with $\alpha\geq 11$. Let $p$ represent the $L_{\infty}$ sensitivity of deduplication, quantization and mean imputation. We also assume the size of the dataset $n\geq 101$ for PCA, and the Lipschitz and smoothness parameters, along with global sensitivity, are set to $1$. We omit the privacy guarantees of deduplication due to space constraint. However, we note that the privacy guarantees for deduplication are the same as those for quantization with $\eta=1$. See Appendix D for details.
Quantization Mean imputation PCA Standard Scaling
${\mathcal{M}}_{G}$ ${1.05\alpha\varepsilon^{2}}\left({1+\eta^{2}p^{2}}\right)$ ${1.05\alpha\varepsilon^{2}}\left({1+\frac{4p^{2}}{(n-p)^{2}}}\right)$ ${1.05\alpha\varepsilon^{2}}\left({1+\frac{12.2^{2}}{(\delta_{\min}^{k})^{2}}}\right)$ ${1.05\alpha\varepsilon^{2}}\left({1+\frac{4}{\sigma_{\min}^{3}}}\right)$
${\mathcal{A}}_{\mathrm{GD}}$ ${1.05\alpha\varepsilon^{2}}\left({4+\eta^{2}p^{2}}\right)$ ${4.2\alpha\varepsilon^{2}}\left({1+\frac{p^{2}}{(n-p)^{2}}}\right)$ ${1.05\alpha\varepsilon^{2}}\left({4+\frac{12.2^{2}}{(\delta_{\min}^{k})^{2}}}\right)$ ${4.2\alpha\varepsilon^{2}}\left({1+\frac{1}{\sigma_{\min}^{3}}}\right)$
${\mathcal{M}}_{L}/{\mathcal{M}}_{E}$ $\varepsilon\left({1+\eta p}\right)$ $\varepsilon\left({1+\frac{2p}{n-p}}\right)$ $\varepsilon\left({1+\frac{12.2}{\delta_{\min}^{k}}}\right)$ $\varepsilon\left({1+\frac{4}{\sigma_{\min}^{3}}}\right)$
${\mathcal{A}}_{\mathrm{SGD-samp}}$ ${1.05\alpha\varepsilon^{2}}\left({2\alpha+\frac{12.2^{2}}{(\delta_{\min}^{k})^{2}}}\right)$ ${2.1\alpha\varepsilon^{2}}\left({\alpha+\frac{8}{\sigma_{\min}^{6}}}\right)$
${\mathcal{A}}_{\mathrm{SGD-iter}}$ ${1.1\alpha\varepsilon^{2}}\left({1+\frac{\frac{\eta^{2}p^{2}n}{\log n}}{\frac{n-p}{\log n-p}}}\right)$ ${1.1\alpha\varepsilon^{2}}\left({1+\frac{4p^{2}\frac{n}{\log n}}{\frac{(n-p)^{3}}{\log n-p}}}\right)$
@@
Figure 2: Visualization of the overall privacy of pre-processed Gaussian mechanism analysed with group privacy and our bound from Theorem 2. Here, $\eta$ is the distance threshold of the quantization algorithm, $n$ is the size of the possible datasets, and $\delta_{\min}$ is the minimum gap between the $k^{th}$ and the $k+1^{th}$ eigenvalue of all possible datasets.
@@
Table 2 shows that the pre-processing methods discussed in Section 4.1 typically lead to a minimal, constant-order increase in the privacy cost for common DP mechanisms, depending on the datasets in ${\mathcal{L}}$. Specifically, deduplication results in a constant increase in the privacy cost when the datasets in ${\mathcal{L}}$ do not contain large clusters, i.e. $\Delta_{\infty}=o(n)$, whereas quantization can handle larger clusters of size $\Delta_{\infty}=o(\nicefrac{{n}}{{\eta}})$. The privacy cost of mean imputation remains constant for datasets with few missing values ($p=o(n)$). Standard scaling leads to constant amplification in privacy cost when all features in dataset collection ${\mathcal{L}}$ have non-vanishing standard deviation. For PCA, we only present the results for rank reduction and the results for dimension reduction follows similarly. The additional privacy cost remains small for low rank datasets with $\delta_{\min}$ bounded away from 0.
@@ 5.1 Limitations of conditional privacy guarantee
It is important to recognise that datasets exhibiting pathologucal characteristics, like those above, are usually not of practical interest in data analysis. For instance, applying mean imputation to a dataset where the number of missing points is nearly equal to the size of the dataset or implementing approximate deduplication that results in the removal of nearly all data are not considered sensible practice. A possible solution to this problem is to potentially refuse to provide an output when the input dataset is deemed pathological, provided that the decision to refuse is made in a manner that preserves privacy. This approach is adopted in the following section, where we employ the Propose-Test-Release framework proposed by Dwork & Lei (2009) to establish unconditional privacy guarantees over all possible datasets, at the expense of accuracy on datasets that are “pathological”.
@@ 5.2 Unconditional privacy guarantees via PTR
We present Algorithm 1 that applies the PTR procedure to combine the non-private PCA for rank reduction with DP-GD. It exhibits unconditional privacy guarantee over all possible datasets (Theorem 8). While this technique is specifically described for combining PCA with DP-GD, it’s worth noting that the same approach can be applied to other pairings of DP mechanisms and pre-processing algorithms, though we do not explicitly detail each combination since their implementations follow the same basic principles.
@@
Input: Dataset $S$, estimated lower bound $\beta$ of $\delta_{k}(S)$, privacy parameters $\varepsilon,\delta$, Lipschitzness $L$
@@
For any $L$-Lipschitz and $\mu$-smooth loss function $\ell$, $\varepsilon>0$ and $\delta\leq\exp\left({-1.05\varepsilon^{2}\left({1+\frac{12.2^{2}\mu^{2}}{L^{2}\beta^{2}}}\right)}\right)$, Algorithm 1 with privacy parameters $\varepsilon,\delta$, and estimated lower bound $\beta$ is $\left({\widehat{\varepsilon}+\varepsilon,\delta}\right)$-DP on a dataset of size $n\geq 101$, where $\widehat{\varepsilon}=3\varepsilon\sqrt{1.05\left({1+\frac{12.2^{2}\mu^{2}}{L^{2}\beta^{2}}}\right)\log\frac{1}{\delta}}$.
@@
For $\delta,\varepsilon$ defined in Algorithm 1, $n\geq 101$, and any $\beta\leq\delta_{k}\left({S}\right)$, with probability at least $1-\xi$,  Algorithm 1 outputs $\hat{\theta}$ such that the excess empirical risk
@@
where $\xi=\frac{1}{2\delta}\exp\left({-\frac{(\delta_{k}(S)-\beta)n\varepsilon}{12.2}}\right),\Lambda=\sum_{i=k+1}^{d}\lambda_{i}(S)$ where the high probability is over the randomness in Step 1 and the expectation is over the randomness of Step 6.
The results follow a similar analysis as Song et al. (2021), which shows that the convergence bound of DP-GD scales with the rank of the dataset. However, when the dataset remains full rank but the first $k$ eigenvalues dominates the rest, e.g. $\Lambda=O(\sqrt{k})$, the convergence bound following Song et al. (2021) is $O(\sqrt{d})$. In contrast, Proposition 9 leads to a dimension-independent convergence bound of order $O(\sqrt{k})$, with a slight degradation in the privacy guarantee. Here, $\beta$ introduces the trade-off between privacy and utility. Large $\beta$ leads to tighter privacy guarantee (small effective $\varepsilon$ in Theorem 8) at the risk of worse utility (large $\xi$ in Proposition 9).
@@
We conducted experiments on a synthetic approximately low rank dataset to corroborate our results in Proposition 9, and summarised the results in Figure 3. We generate a 2-class low rank dataset consisting of 1000 data points with dimension 6000 and approximately rank 50. The synthetic dataset has positive yet small eigenvalues for the $k^{\text{th}}$ eigenvectors for $k\geq 50$, ensuring $2L\Lambda$ in Equation 4 is small but positive.
For each overall privacy parameter $\varepsilon$, we evaluated the excess empirical risk of DP-SGD with: a) non-private PCA with an adjusted privacy parameter from Table 2, b) DP-PCA, with half of the privacy budget to allocated to DP-PCA, and c) no pre-processing.
Figure 3 shows that pre-processed DP pipeline outperforms logistic regression without pre-processing. This is because the dataset’s approximate low rankness, indicated by $02L\Lambda>$ 0, prevents logistic regression from leveraging the dimension-independent optimization guarantees using the original high dimensional datasets (Song et al., 2021). However, non-private PCA, while incurring a minor constant-order privacy cost, effectively utilizes the data’s low rankness and offers significant benefits, especially at smaller $\varepsilon$ values where optimization error is dominated by the first term in Equation 4Figure 3 also demonstrates that DP-PCA exhibits the worst performance among the three methods, beccause the inherent error of order $\Omega(d)$ of DP-PCA dominates given a high dimensional dataset with $d\ll n$(Liu et al., 2022).
In Appendix F, we provide the details on experimental setups and a discussion on the practical modification of Algorithm 1 with clipping.
@@ 6 Conclusion
Acknowledgements
We thank Francesco Pinto for the initial discussion of the project. We also thank Alexandru Ţifea, Jiduan Wu and Omri Ben-Dov for their useful feedback. Additionally, we extend our gratitude to the reviewers of ICML for their careful reviews and insightful suggestions.
@@ Acknowledgements
In this section, we present the proofs regarding SRDP properties. In Section A.1, we introduce SRDP satisfies some desired properties, and present their proofs. Section A.2 provides the exact SRDP parameters of common DP mechanisms.
A.1 Properties of SRDP and their proofs
While conditional SRDP over a set ${\mathcal{L}}$ can be considered as a special case of Pufferfish Rényi Privacy (Kifer & Machanavajjhala, 2014), we show that it satisfies desirable properties, such as sequential composition and privacy amplification by subsampling, which are not satisfied by the more general Pufferfish Renyi Privacy (Pierquin & Bellet, 2023).
Properties of SRDP
Lemma 2.
Let ${\mathcal{L}}$ be any dataset collection. Then, the following holds.
See 2
@@ A.1 Properties of SRDP and their proofs
Lemma A (Additivity of Renyi divergence (van Erven & Harremoës, 2012)).
@@
Lemma B (Weak convexity of Renyi divergence, Lemma 25 in Feldman et al. (2018)).
@@ A.2 Proof of Theorem 1(Derivation of Table 1)
A loss function $\ell:\Theta\times{\mathcal{K}}\to\mathbb{R}_{+}$ is $\mu$-smooth if for all $\theta\in\Theta$ and $x,x^{\prime}\in{\mathcal{K}}$, $\left\lVert\nabla_{\theta}\ell(\theta,x)-\nabla_{\theta}\ell(\theta,x)\right\rVert_{2}\leq\mu\left\lVert x-x^{\prime}\right\rVert_{2}$.
We note that Definition 6 is different from the usual definition of smoothness in the literature that requires $\left\lVert\nabla_{\theta}\ell(\theta,x)-\nabla_{\theta}\ell(\theta^{\prime},x)\right\rVert_{2}\leq\mu\left\lVert\theta-\theta^{\prime}\right\rVert_{2}$ for all $x\in{\mathcal{K}}$ and $\theta,\theta^{\prime}\in\Theta$ (Chaudhuri et al., 2011; Feldman et al., 2018). However, many loss functions, including square loss and logistic loss, are smooth on common model classes based on Definition 6.
@@ A.2.1 Output Perturbation and Random Sampling methods
This section provides the detailed theorems and proofs for Smooth RDP parameters of output perturbation methods (Gaussian and Laplace mechanism) and random sampling methods (Exponential mechanism) in Theorem 10, 11 and 12 respectively, as stated in Table 1.
@@
Lemma C (Corrolary 3 in Mironov (2017)).
@@
Laplace mechanism ${\mathcal{M}}_{L}$ is $(\varepsilon,0)$-DP (Dwork et al., 2006). By the equivalence of $(\varepsilon,0)$-DP and $(\infty,\varepsilon)$-RDP (Mironov, 2017), ${\mathcal{M}}_{L}$ is $(\infty,\varepsilon)$-RDP. This implies that ${\mathcal{M}}_{L}$ is $(\alpha,\varepsilon)$-RDP for any $1\leq\alpha\leq\infty$ by the monotonicity of Rényi divergence (D).
Lemma D (Monotonicity of Rényi divergence Mironov (2017)).
@@
The proof for RDP of Exponential mechanism is similar to that of Laplace mechanism. As the exponential mechanism ${\mathcal{M}}_{E}$ is $(\varepsilon,0)$-DP (Dwork et al., 2006). By the equivalence of $(\varepsilon,0)$-DP and $(\infty,\varepsilon)$-RDP (Mironov, 2017), ${\mathcal{M}}_{E}$ is $(\infty,\varepsilon)$-RDP. By the monotonicity of Rényi divergence (D), ${\mathcal{M}}_{E}$ is also $(\alpha,\varepsilon)$-RDP for any $\alpha\geq 1$.